For an interest rate of 13 percent compounded annually, how much will be required in
four years to repay a $25,000 loan received now?
1
Expert's answer
2020-10-16T07:26:29-0400

Principal, "P=\$25,000"

"Rate, r= 13\%"

"Time, t=4years"

"n=" 1 compounded annually

Amount to be paid in four years (compounded annually) will be

"A=P(1+\frac{r}{n})^{nt}=25,000" "(1+\frac{13}{100\times1})^{(1)(4)}"

"=25,000(1+0.13)^{4}=25,000(1.13)^{4}=\$40761.84"